OpenStudy (anonymous):

Convert the following polar representation of this complex number into it's regular form: z=4(cos pi/2+isinpi/2)? A.(0,4) B.(-4,0) C.(4sqrt2,0) D.(0,-4)

OpenStudy (jtvatsim):

The question seems a bit strange as currently worded... The answer choices don't match up... since the complex number is ALREADY in polar representation... :)

OpenStudy (anonymous):

ill double check the question

OpenStudy (anonymous):

@jtvatsim fixed it:)

OpenStudy (jtvatsim):

That makes more sense... but now let's see if I can figure out the real question. :P lol

OpenStudy (jtvatsim):

Alright, so for these you need to either have memorized what cos(pi/2) and sin(pi/2) are, use a calculator (in radians mode), or use the unit circle. :)

OpenStudy (jtvatsim):

I happen to see these a lot, so I can tell you that cos(pi/2) = 0 and sin(pi/2) = 1.

OpenStudy (jtvatsim):

OK so far?

OpenStudy (anonymous):

yeah

OpenStudy (jtvatsim):

Alright, so now we just simplify the given number: z = 4(cos pi/2 + i sin pi/2) = 4(0 + i * 1) = 4( i ) = 4i.

OpenStudy (jtvatsim):

If you are feeling really "textbook mathy" this is the same as 4i = 0 + 4i or some people write it like a coordinate (0, 4) where it is understood that the first coordinate gives the "real" number values, and the second coordinate gives the "imaginary" i values.

OpenStudy (anonymous):

so in order to get the answer that's all you do?

OpenStudy (jtvatsim):

Yes, converting from polar form to regular form is much easier than going the other direction.

OpenStudy (anonymous):

Thanks:) So the final amswer is A. (0,4

OpenStudy (jtvatsim):

No problem! You are good to go!